OpenStudy (anonymous):

Hi, I am currently on problem 1F-2 of the first problem set and am having trouble following the algebra. I have reached to this part: 10x^9(x^2 + 1)^10 + x^10[10(x2 + 1)^9(2x)]. I do not understand how they arrive at = 10(3x^2 + 1)x^9(x^2 + 1)^9. Could someone please explain?

OpenStudy (phi):

\[ 10x^9(x^2 + 1)^{10} + x^{10}\cdot 10(x^2 + 1)^9(2x) \] factor out \( 10x^9 (x^2+1)^9 \) from each term: \[ 10x^9 (x^2+1)^9\left( (x^2 + 1)+x\cdot 2x\right) \\ 10x^9 (x^2+1)^9\left( x^2 + 1+2x^2\right) \\ 10x^9 (x^2+1)^9(3x^2+1) \]